### v1

# InfoLossQA: Characterizing and Recovering Information Loss in Text Simplification

Abstract
Text simplification aims to make technical texts more accessible to laypeople but often results in deletion of information and vagueness. This work proposes InfoLossQA, a framework to characterize and recover simplification-induced information loss in form of question-and-answer (QA) pairs. Building on the theory of Question Under Discussion, the QA pairs are designed to help readers deepen their knowledge of a text. We conduct a range of experiments with this framework. First, we collect a dataset of 1,000 linguist-curated QA pairs derived from 104 LLM simplifications of scientific abstracts of medical studies. Our analyses of this data reveal that information loss occurs frequently, and that the QA pairs give a high-level overview of what information was lost. Second, we devise two methods for this task: end-to-end prompting of open-source and commercial language models, and a natural language inference pipeline. With a novel evaluation framework considering the correctness of QA pairs and their linguistic suitability, our expert evaluation reveals that models struggle to reliably identify information loss and applying similar standards as humans at what constitutes information loss.
InfoLossQA: Characterizing and Recovering Information Loss in
Text Simplification
[…]
1 Introduction
Technical texts, many of which exist in high-stake domains (e.g., medicine), are often written in a language incomprehensible to lay readers. Improving the accessibility of texts may help address wider social issues such as disinformation UNSDG (2021) and improve access to higher education Goff et al. (2004). Automatic text simplification that rewrites text into plain language may therefore be a technology for good. With the adoption of LLMs, document-level text simplification has significantly progressed in recent years August et al. (2023); Laban et al. (2023); Agrawal and Carpuat (2023).
Simplification is an inherently lossy process: Even when done by professional editors, the resulting plain language tends to lose details and become more generic Li and Nenkova (2015), and some content is omitted Zhong et al. (2020). The over-simplification of content—including excessive deletion and vagueness—may lead to misinterpretations and factual errors Devaraj et al. (2022), or reduced comprehension Agrawal and Carpuat (2023). Therefore, we consider the following question: How can we characterize information loss and help readers recover what is lost in an intuitive and understandable manner?
Figure 1: The goal of InfoLossQA is to generate a series of QA pairs that reveal to lay readers what information a simplified text lacks compared to its original.
This paper presents a question generation and answering framework to characterize and recover information loss during simplification in a reader-centric manner. Inspired by the theoretical work of “potential questions” Onea (2016) and expectation-driven Questions Under Discussion (QUD, Kehler and Rohde, 2017), information loss is characterized in the form of pragmatically felicitous questions rooted in simplified texts, without assuming access to the original text. Information loss is then recovered by answering these questions in plain language, based on the original (technical) text (Figure 1). Informed by QUD, the questions are curiosity-driven and information-seeking, and are subject to well-defined linguistic constraints Riester et al. (2018); Wu et al. (2023). While prior work used factoid QAs to describe specific factual changes Cole et al. (2023), the QUD framework suggests higher-level questions, and allows us to characterize general yet subtle language changes pervasive in factually equivalent texts. These include lower specificity Louis and Nenkova (2011) and greater degrees of vagueness van Deemter (2012). We envision the resulting QA pairs being directly incorporated in interactive simplification tools Fok et al. (2023).
We first introduce InfoLossQA, a linguist-curated dataset grounded in the scenario of lay readers digesting simplified abstracts of randomized controlled trials (RCTs) in medicine and healthcare. We focus on RCTs because they combine several important and illustrative properties: They use highly technical/inaccessible language; lay audiences often having keen interest in their content; and misinterpretations are risky, given the implications for health. InfoLossQA consists of 1,000 QA pairs across 104 RCT abstracts simplified by GPT-4 (prior work has shown that GPT-3.5 is a capable simplifier of medical texts; Shaib et al., 2023). The QA pairs provide a rich description of what information is lost both on a pragmatic and semantic level. Our extensive analysis shows that technical sections of RCTs are particularly prone for information loss. Questions often elaborate about why and how a clinical trial was conducted, going beyond frequently asked questions previously identified in this domain August et al. (2023).
We then present two methods to perform the question generation and answering task. First, an end-to-end model, where we prompt open-source and API-based LLMs with detailed task instructions and a one-shot example. Second, a natural language inference (NLI) pipeline which reasons over entailment of atomic facts Stacey et al. (2023).
Finally, we design a comprehensive framework to evaluate models at this task. The framework includes 10 quality desiderata evaluating the correctness of QAs, their linguistic suitability, and the recall of human-written QAs. Expert judgments are collected on more than 400 generated QA pairs. We find that models have good question generation and answering capabilities but fall short at reliably identifying information loss. In this regard, the NLI pipeline is more effective than open-source LLMs, but produces QA pairs with a smaller granularity.
To summarize, we make the following technical contributions:
  1. 1.
    We propose InfoLossQA, a novel task and linguist-curated dataset to characterize and reveal simplification-induced information loss in form of QA pairs.
  2. 2.
    We establish two baselines at this task, including end-to-end LLMs and an NLI pipeline.
  3. 3.
    We propose a comprehensive framework to evaluate models at this task and collect human quality judgments.
We released the experiment code and the dataset at https://InfoLossQA.ikim.nrw/.
2 The InfoLossQA Task
Information loss is inherent to simplification, insofar as the task demands producing outputs which are both comparatively uncomplicated and reasonably concise. Characterizing and revealing the specific content elided is critical to (1) provide a path for users to achieve better understanding than only seeing the simplified text, supporting users’ growth Gooding (2022), and (2) allow researchers to evaluate the quality of a simplified text, given that omitting or oversimplifying key information may yield misleading texts Devaraj et al. (2022). Moreover, we envision simplification as an interactive process in which a user may ask for additional details on particular points. InfoLossQA defines a task and dataset that entails generating QA pairs revealing to lay readers the information missing from a simplified text, compared to its original.
Task formulation.
Given a pair of documents, the original text $X_{\text{src}}$ and the simplified text $X_{\text{tgt}}$, the task entails generating QA pairs that each reveal to lay readers one instance of information loss. We define an instance of information loss as a tuple $(X_{\text{src}},X_{\text{tgt}},Q,A,E,R)$. A potential question $Q$ asks for information absent from $X_{\text{tgt}}$. The answer $A$ to the question $Q$ provides the missing information. We additionally require two forms of localization or grounding to support richer analysis and scrutiny of errors in the generated QA pairs. First, each answer $A$ must be supported by evidence $E\in X_{\text{src}}$ extracted from the original text. Second, each $Q$ and $A$ pair may be supported by an extractive rationale $R\in X_{\text{tgt}}$ which localizes where the information loss or vagueness occurs within the simplified text. Note that while $E$ always exists for an instance, $R$ may not (some information in $X_{\text{src}}$ may be completely absent from $X_{\text{tgt}}$). This formulation permits multiple information loss instances for a given $X_{\text{src}}$ and $X_{\text{tgt}}$ pair. We do not constrain the length or linguistic unit(s) of the extractive information loss spans ($E$ and $R$): they can comprise words, phrases, sentences, or entire paragraphs, and a single $E$ or $R$ can be one or more spans.
Types of information loss.
For the purposes of this study, we define two types of information loss:
  1. 1.
    Deletion. Pieces of information which were not included in the simplification.
  2. 2.
    Oversimplification. Pieces of information that were simplified to the extent that they are vague or devoid of their original meaning. This is the case where the extractive rationale $R$ is identified.
These categories are meant to be flexible to capture a variety of information transformations, but also sufficiently concrete to be operationalized in an annotation protocol (Section 3).
Linguistic suitability.
One of our primary goals is to generate QA pairs that might help users gain a fuller understanding of the original text Fok et al. (2023). The desire for a deeper understanding may arise when a user has a keen interest in aspects of the original text that go beyond what is presented in the simplified version. Enabling such interactions may also allow individuals to calibrate their trust in AI-simplified outputs McDermott and Brink (2019). We pose two requirements for the QA pairs:
(1) Readability. The language level of the questions and answers should match that of the simplified text. That is, while the QA must discuss technical material from the original text, it should be explained at a level appropriate for the reader. [^1] It is possible that $X_{\text{tgt}}$ itself is not at the right level for a given reader. However, this is out of scope of our study.
(2) Givenness. Questions should be pragmatically felicitous to be understood by a reader without having seen the answer or the original text. The theory of QUD formalizes this through the Givenness constraint Riester et al. (2018); Wu et al. (2023) which specifies that $Q$ should not contain concepts that are hearer-new Markert et al. (2012) with respect to a common ground. Here, the common ground is the simplified text $X_{\text{tgt}}$. Intuitively, this means that it should be clear from reading the question how the answer would expand on what a reader already knows from the simplified text.
Example.
We illustrate the different elements and linguistic constraints of the task with an example. Consider the following passages taken from the abstract of a clinical study:
Original ($X_{\text{src}}$): These results indicate that acute/chronic endurance ($E_{1}$) arm-cranking with EMS applied to the lower limbs improves the brachial artery endothelial function ($E_{2}$) more markedly than the same exercise without EMS.
Simplified ($X_{\text{tgt}}$): The study concluded that doing the arm-cranking exercise with EMS on the lower body can improve arm function ($R_{2}$) more than doing the exercise without EMS.
This example demonstrates the challenging nature of the task. To identify information loss, models need to bridge two language levels to decide on the presence and nature of information loss. Relying on lexical overlap, ordering, or other surface level properties is infeasible. The example includes both a deletion (EMS helps in “acute/chronic” settings) and an oversimplification (“improve arm function” is too broad given that EMS improves “artery function”). By contrast, the lexical simplification “lower limbs” $\rightarrow$ “lower body” may be viewed as functionally equivalent, and an arguably unproblematic case of information loss.
This simplification gives rise to two QA pairs:
  1. Q1.
    On what timeframes does EMS improve training? A. It helps in single application (acute) or repeated application (chronic).
  2. Q2.
    How did the researchers measure how well EMS works? A. The researchers measure to what extent the main artery of the arm widens, which is called flow-mediated vasodilation (FMD).
Note that the QA pairs fulfill the stylistic and linguistic constraints we defined above: the technical material is explained as necessary, and the questions are interpretable for readers without having seen the original text fulfilling the givenness constraint. For contrast, an invalid (↯) question conflicting with the givenness constraint is as follows:
  1. Q3.
    Did EMS improve training in acute and chronic applications?
While Q1 and Q3 ask about the same information, Q1 would be more natural to lay readers because it could conceivably be asked without having seen the original text (it does not presuppose that the study looked at short-term and long-term effects).
3 Data Collection
3.1 Simplification Corpus
In this work, we focus on medical texts, which provide an important and illustrative case where lay individuals may be keenly interested in understanding newly published science. More specifically, we collect annotations on abstracts of medical journal articles that describe the conduct and findings of Randomized Controlled Trials (RCTs). RCTs are the foundation of evidence-based medicine Sackett (1998) and provide the critical comparative effectiveness findings that inform patient care. But articles describing RCTs use technical language, and so are effectively inaccessible to (lay) patients Zuccala (2010); August et al. (2023). This means that individuals may already interact with LLMs to simplify such texts. But automated simplification of RCTs may result in information loss, which in turn risks consumers misinterpreting findings.
To better quantify information loss in LLM-simplified outputs, we create a parallel simplification corpus with GPT-4 (gpt-4-0613). [^2] There is no existing curated parallel corpus for RCT simplification. Initially, we considered obtaining medical abstracts that are not RCTs with their associated Plain Language Summaries (PLS) from journals, but found that these PLS vary too significantly in content and style to form a parallel corpus, a problem also noted in prior related work Devaraj et al. (2021), partly due to the ongoing debate what an effective PLS should look like Jelicic Kadic et al. (2016). On the related task of plain language summarization of full RCT publications, Shaib et al. (2023) observed very low error rates (less than 10% of summaries affected), but we do not yet have a quantitative understanding of information loss in LLM-simplified RCT texts. We sample RCT abstracts from the Evidence Inference dataset DeYoung et al. (2020b); Lehman et al. (2019), originally sourced from PubMed. We provide complete generation parameters and example outputs in Appendix A.
3.2 Annotation Process
The process for annotating InfoLossQA is as follows. First, annotators examine both the original and simplified versions of the text. Next they identify and highlight any spans where information is missing from the simplified text. For each highlighted instance, annotators compose a QA-pair.
(a) Distribution of QA pairs over RCT abstract sections. Some QAs are localized in multiple sections (multi.) or in a rare section (other). We also report the percentage of QAs that falls into abstracts without section structure (19.2% of abstracts; no secs.).
(b) Number of QA pairs per sentence for a random selection of 25 documents. The x-axis are sentence indices.
Figure 2: Distribution of information loss over sections and sentences. Humans produce a similar distribution of questions by section (a), but the questions differ in their localization (b). Comparing humans to models, we can see significant differences where questions are asked, and by extension also what they are about.
Scenario grounding.
Technical texts like RCTs have a long tail of information that can be lost during simplification. For example, numerical results like confidence intervals and significance scores. Annotating all information loss is intractable, and the resulting QAs would likely overwhelm readers. Therefore, we grounded annotation in a scenario: annotators were asked to imagine a patient who could benefit from better understanding the RCT at hand, and instructed to focus on information that may be salient in this scenario.
This notion of saliency is subjective as people have different information needs. Similarly, prior work in QUD Ko et al. (2022) showed that question generation by human annotators is also subjective: People ask different questions even when the answer to these questions is the same. In our dataset, each document was annotated by two annotators, independently. This allows us to evaluate the extent to which people ask similar questions (Section 3.3).
Annotators.
We hired three expert annotators majoring in linguistics with experience annotating medical texts. To ensure high annotation quality, we conducted a pilot phase with written feedback and held frequent review meetings to discuss questions. Annotators were paid \$15/hr. [^3] Higher than local minimum wage at the time of study. The median duration per document was 22 minutes. We used the Thresh toolkit Heineman et al. (2023) for annotation which allows parallel text highlights. See Appendix B for annotation guidelines and screenshots of the interface that we used.
Dataset statistics.
We collected annotations on 104 abstract-simplification pairs (average number of tokens: $|X_{\text{src}}|=312$; $|X_{\text{tgt}}|=271$) yielding 1,000 questions and answers (average number of tokens: $|Q|=11.4$; $|A|=30.3$). [^4] Tokenizing with nltk Bird et al. (2009). Each annotator wrote on average between 4.1 and 5.7 QA pairs (i.e., information loss instances) per document. We observe that oversimplification is more prevalent (74.1% of questions) in this corpus than complete deletion of information (25.9% of questions). A detailed breakdown of the statistics by annotator is given in Table A.2 and Section 3.3 analyzes to what extent annotators ask similar questions.
Question Type % All % O % D Examples
Procedural. Asking for the procedures, tools, or methods by which a certain outcome is achieved.
34.3 34.1 34.7
• How did they measure the patients’ opioid medication needs?
• How did the study control for bias?
Concept. Asking for a definition of an event or a concept.
25.7 28.1 18.9
• What kind of hip surgery were patients undergoing? • What type of mental illnesses are being studied?
Extent. Asking for the extent or quantity of an event or a concept.
17.0 19.2 10.8
• How many patients were in each group? • How much lurasidone was given to the patients in the trial?
Comparison. Asking for comparison among multiple events or concepts.
8.3 8.9 6.6
• How much did the abnormal blood vessels reduce in group 1 compared to group 2? • How did pre-meal insulin affect platelet activation compared to the placebo?
Consequence. Asking for the consequences or results of an event.
5.7 3.0 13.5
• What else does increased blood flow do? • What was the correlation between measures for total daily calcium intake?
Cause. Asking for the cause or reason for an event or a concept.
4.3 1.6 12.0
• What was a motivation behind the study? • Why are soy products being looked at in this study?
Example. Asking for example(s) or instance(s) of an event or a concept.
3.4 4.3 0.8
• What are the other brain functions that were tested? • What medication is typically used for outpatients?
Disjunction. Asking for the true one given multiple events or concepts, where comparison among options is not needed.
0.6 0.5 0.8
• Which one of the three components did this group’s supplement contain? • Which gel saw the increase in beneficial microbes after 24 hours?
Verification. Asking for the truthfulness of an event or a concept.
0.7 0.3 1.9
• Were there any other changes in blood sugar level observed in the study? • Were both eyes of each patient tested?
Judgmental. Asking for the opinions of the answerer’s own.
Not observed.
Table 1: Example questions in InfoLossQA classified according to the question taxonomy by Cao and Wang (2021). Percent of all questions is given by category (O = Oversimplification; D = Deletion).
3.3 Analysis of InfoLossQA
We analyze the fine-grained annotation of questions and their localization in form of evidence spans to get a better understanding of how models preserve (or omit) content during simplification.
Where in the text is most information lost?
We combine the precise localization of a QA pair with structural elements of the abstracts to identify areas where the most information loss occurs. By convention, most RCT abstracts are structured, i.e., having section headings for Background, Methods, Results and Conclusion (most commonly). We use a regular expression to identify the start and end boundaries of sections. Because the specific naming of such sections differs across articles, we collected a lookup dictionary to normalize synonymous headers. [^5] For example, “Design,” “Methods” and “Materials and Methods” are all mapped to “Methods.” Afterwards, we align QA pairs to sections based on their localization.
Figure 2 shows the distribution of QAs over RCT sections. Most information loss occurs in the Methods section (41.7% of QAs). This is intuitive as Methods sections are likely to be particularly technical. The least information loss occurs in the Conclusion (3.5% of QAs) section. A small percentage of QAs spans multiple sections. These tend to be about concepts that were repeatedly oversimplified.
What types of questions are asked?
To characterize the information typically lost, we categorize questions according to 10 pragmatic function types defined by Cao and Wang (2021). To automate this analysis, we designed a few-shot prompting classifier based on the annotation guidelines of the authors. [^6] Prompting gpt-4-0613 with few-shot in-domain examples achieves an F1 score of 0.87 on a validation set. See Appendix A for details. Table 1 shows the question distribution across types, their definition and examples.
Most questions are of type procedural, requesting specific details about how the study was conducted (34.3%). This aligns with the earlier observation that most information loss occurs in Methods sections. The second most common type of questions seek more concrete definitions for overgeneralized concepts (25.7%), such as the health condition or clinical procedure. Comparing deletions and oversimplifications, we observe a similar distribution, with the exception of questions asking for the consequences (e.g., experimental results) or causes (e.g., what motivated the study) of an event being more prevalent among deletions.
In addition to the domain-agnostic question taxonomy above, we also analyze the QA pairs through the lens of the Paper Plain Key Question Index which was designed to convey the most important elements of an RCT August et al. (2023). We manually classify 120 QA pairs (40 per annotator) according to the 8 categories of the question index. Aligned with the above findings, the information which is most likely lost are about methodological detail (i.e., study protocol, analysis tools, population; 53% of QA pairs), followed by results (18% of QA pairs) and the goals of the trial (11% of QA pairs). The most critical information (usual and new treatment) are preserved. See Table A.4 for a detailed breakdown and examples.
How similar are information needs across annotators?
As discussed in Section 3.2, the InfoLossQA task is subjective. Enlisting two annotators per document allows us to quantitatively study information needs of different individuals.
First, we consider the extent to which annotators highlight the same information loss spans. We calculate three measures on the evidence spans $E$ of the original text: (1) strict span-level F1, where boundaries and type (deletion vs. oversimplification) have to match exactly, (2) partial span-level F1 where two spans are considered a match if their token-level intersection over union exceeds a threshold DeYoung et al. (2020a), [^7] We tokenize on whitespace and use a threshold of $t=0.5$. and (3) sentence-level Krippendorff’s $\alpha$ where we project span-level annotations to a binary label indicating if a sentence has at least one span Goyal et al. (2022). As shown in Table 2, the sentence-level $\alpha$ shows that the annotators do tend to pick up information loss localized in the same evidence sentence to a moderate degree Artstein and Poesio (2008). However, the precise evidence span (in smaller linguistic units) is more subjective, as indicated by the lower span-level scores.
Second, we analyze the similarity of questions $Q$ as a function of whether their underlying evidence spans E have complete, partial, or no overlap. Generally, questions become more similar when the underlying evidence is the same (Figure 3). However, subjectivity remains: Some questions are different even though the evidence is the same. Similar observations were made for human-generated QUDs Ko et al. (2022).
Granularity (Measure) Del. Over. All
Span, exact (F1) 0.20 0.13 0.17
Span, partial (F1) 0.34 0.34 0.34
Sentence (Krippendorff’s $\alpha$) 0.50
Table 2: Evidence span ($E$) overlap between annotators. On a span-level, overlap is reported separately for deletions (del) and oversimplifications (over).
Figure 3: Similarity of annotators’ questions ($Q$) by whether their evidence spans ($E$) have no/partial/complete overlap. We calculate cosine similarity over sentence embeddings Reimers and Gurevych (2019). Cosine greater 1 is an artifact of the kernel density estimate.
4 Methods for Automatically Identifying Information Loss
No system in existing literature is directly applicable to solve all parts of the InfoLossQA task. We propose and evaluate two systems: (1) End-to-end LLM prompting, contrasting several strong open-source and commercial LLMs; (2) A pipeline that combines ideas from fact extraction, NLI, and question-generation into a full system.
4.1 LLM Prompting
We design a one-shot prompt that contains a detailed specification of the task, closely resembling the manual annotation guidelines. [^8] Compared to a zero-shot prompt, we found that even one example substantially improves performance. For a fair comparison across models, we do not include more than one example as they would exhaust the context windows of some models under investigation. We ask all models to localize information loss instances, provide a short motivation for why this instance is considered information loss, akin to Chain-of-Thought prompting Wei et al. (2022), and finally to generate the QA pairs. We benchmark three LLMs: Mistral-7B, Llama2-70B, and GPT-4.999Specifically, we use gpt-4-0613, llama-2-70b-chat, and mistral-7b-instruct-v0.1. All prompts were manually tuned for each model, though this was a heuristic process and better prompts may exist. See Appendix A for prompts and decoding parameters.
4.2 NLI Pipeline
To contrast with end-to-end LLMs, we implement a three-step pipeline: (i) fact extraction, (ii) entailment prediction, (iii) QA generation for facts with a neutral relation.
NLI aims to determine whether a “hypothesis” can be inferred from a “premise,” classifying the hypothesis as entailed, contradictory or neutral. We use this framework to reason about the entailment of information in the original text (= hypothesis) with respect to its simplification (= premise). Intuitively, a neutral relation should hold when there is at least one piece of information in the original text which cannot be inferred from its simplification, i.e., an instance of information loss. In contrast, entailment indicates that all information is conveyed, and contradictions should not occur if the simplification is faithful.
A key challenge is the granularity of relation classification. The predominant approach in NLI is to classify at the sentence-level Bowman et al. (2015) with some extensions to document-level Schuster et al. (2022); Yin et al. (2021). However, for our purposes sentence-level is too coarse. Information loss instances can be as fine-grained as a single word or phrase, and one sentence may contain multiple distinct instances. To address this, we follow Stacey et al. (2023), who proposed entailment classification at the level of atomic facts. Reasoning over fine-grained facts is an established strategy, for example, in summarization evaluation Nenkova and Passonneau (2004); Liu et al. (2023), but a bottleneck is obtaining such facts. Stacey et al. (2023) showed that LLMs can effectively decompose sentences into shorter sentences, each encapsulating an atomic fact.
We reconcile these ideas into a pipeline as follows. First, we use an LLM to decompose each sentence in the original text $X_{\text{src}}$ into a list of atomic facts. Then, a standard NLI model Laurer et al. (2024) classifies the entailment of each fact against the simple text $X_{\text{tgt}}$. Finally, we prompt an LLM to generate QA pairs given $X_{\text{src}}$, $X_{\text{tgt}}$, and the list of neutral/missing facts. The sentence that generated the fact serves as evidence $E$ [^10] We do not implement localization $R$ in the simple text, but note that interpretable NLI models could provide such a localization Stacey et al. (2022); Rodriguez et al. (2023). See Appendix A for prompts and decoding parameters.
5 Evaluation Framework
We next propose a comprehensive framework to evaluate automatic systems for InfoLossQA. This framework considers: (i) Correctness of the questions and answers (i.e., the QA pairs); (ii) Linguistic suitability for user-facing settings, and; (iii) Coverage of human QAs. All evaluation is done manually by trained annotators. We outline the evaluation criteria below and provide full annotator guidelines and examples in Appendix C.
5.1 Evaluation Criteria
Angle 1: Correctness of generated QA pairs.
Our first criterion, Answerability (3-point), states that the generated questions ($Q$s) should be about an information loss. Valid questions must be answerable with the original text (i.e., there exists an evidence $E$ answering $Q$), but unanswerable or only vaguely answerable with the simple text.
The second criterion, Answer Accuracy, states that questions should be correctly answered by the generated/extracted answers ($A$s). We evaluate this in three parts. First, Accuracy - Evidence (3-point) captures how well the highlighted evidence $E$ answers $Q$. Second, Accuracy - Answer (3-point) evaluates how well $A$ answers $Q$. Finally, we ask if $A$ contains any Hallucinations (binary).
Angle 2: Linguistic suitability.
As discussed in Section 2, we aim for QA pairs that are pragmatically felicitous, and written in a style that is comprehensible for readers. We propose evaluating this through four linguistic criteria. First, we evaluate Givenness (binary), closely following the constraints of QUD Riester et al. (2018); Wu et al. (2023): A question should only contain concepts that are either directly mentioned in the question context, inferable, or generally known. We define the question context as the entire simple text $X_{\text{tgt}}$ for QAs on Deletions, and everything up to and including the question localization $R$ for QAs on Oversimplifications. [^11] With this definition of question context we intend to enable interactions where readers can choose to inspect Oversimplification QAs as they read the text, for example, visualized as expandable highlights Fok et al. (2023).
Second, when a question seeks clarification about an oversimplified concept, there should be a rationale span $R$ in the simplified text. We evaluate this with the Rationale Localization criterion (3-point: good, unrelated, missing). Third, we evaluate if $A$ contains any Jargon (4-point). Finally, answers should be Standalone (binary): they should not contain any unresolved coreferences or acronyms that would require reading the original text. This is akin to the requirements for decontextualization Choi et al. (2021).
Angle 3: Recall of human QA.
The third angle evaluates if models identify similar information loss as humans, i.e., the recall of reference QAs in InfoLossQA. The challenge in this evaluation is that both humans and models produce a set of QAs without a strict 1-to-1 correspondence. It is possible that a QA on one side semantically entails zero, one, or more QAs on the other side. In a similar setting, Chen et al. (2022) directly compare the two sets, asking annotators to mark all QA pairs which are not entailed by the other set. However, we found this setup too cognitively demanding with typical sets of more than five QAs.
Therefore, we opt for a pairwise setting, comparing each reference QA with one model QA at a time, and marking the reference as either fully recalled, partially recalled or not recalled. As comparing all possible pairs is intractable, we follow a paraphrase mining approach to get candidate matches Wright et al. (2022). We use sentence embeddings Reimers and Gurevych (2019) to determine the similarity of two QA pairs. Afterwards, we establish a threshold by annotating 400 pairs equally sampled from 20 bins in the similarity range of $[0,1]$ and observe the first bin with a notable number of recalled QA (here: $T=0.65$). Pairs with similarity $\leq T$ are set to not recalled. [^12] For our dataset of 1,000 reference QAs across 104 documents and predictions by four models, this process reduced the total necessary comparisons by 93% from 33,825 to 2,446.
5.2 Annotation Protocol
We sample 100 QA pairs per model for angle 1+2, and 50% of the documents for angle 3.131313For angle 1+2, we take a stratified sample to preserve the relative frequencies at which models generate QAs per RCT section. As the generations by Mistral-7B and Llama2-70B do not always follow the specified output format, we only sample from QA pairs which could be completely parsed. See Table A.3 for a detailed analysis of generation errors. All samples were evaluated by two experts from the pool described in Section 3.2. The total evaluation effort was 64 hours for angle 1+2, and 20 hours for angle 3, with additional training hours. We observe a generally high agreement with slightly lower scores for Jargon and Answerability (Table 3).
Criterion Kappa
$Q$ is Answerable w/ $X_{\text{src}}$ 0.61
$Q$ is Answerable w/ $X_{\text{tgt}}$ 0.50
Accuracy - Evidence ($E$) 0.71
Accuracy - Answer ($A$) 0.68
Hallucinations ($A$) 0.89
Givenness ($Q$) 0.79
Rationale Localization ($R$) 0.80
Jargon ($A$) 0.57
Standalone ($A$) 0.69
Recall of human QA 0.70
Table 3: Inter-annotator agreement for manual evaluation: Randolph’s Kappa Randolph (2005).
$Q$ is Answerable w/ $X_{\text{src}}$ $Q$ is Answerable w/ $X_{\text{tgt}}$ Accuracy - Evidence ($E$) Accuracy - Answer ($A$)
Model
Yes
${}^{\downarrow}$
${}^{\downarrow}$
No
Vague
${}^{\downarrow}$
Correct
${}^{\downarrow}$
${}^{\downarrow}$
Correct
${}^{\downarrow}$
${}^{\downarrow}$
Mistral-7B
72.5
19.6
7.8
32.4
30.4
37.3
74.5
12.7
12.7
84.8
12.3
2.9
Llama2-70B
83.0
9.7
7.3
21.4
32.5
46.1
77.7
11.7
10.7
85.9
10.7
3.4
GPT-4
87.5
12.5
0.0 58.5 33.0 8.5 98.0 2.0 0.0 94.0 6.0 0.0
NLI Pipeline 89.7 7.8
2.5
49.5
25.0
25.5
77.5
4.4
18.1
77.9
20.6
1.5
Table 4: Correctness of generated QA pairs (Angle 1) as percent of ratings given by two independent annotators over 100 QA pairs per model. ${}^{\downarrow}$ = lower is better; Multi. = Multiple valid answers; Part. = Partially complete answer.
Givenness ($Q$) Rationale Localization ($R$) Jargon ($A$) Standalone ($A$) Hallu. ($A$)
Model % Good Good ${}^{\downarrow}$ ${}^{\downarrow}$ (++) (+) (${}^{\downarrow}$ (${}^{\downarrow}$ Good % Good
Mistral-7B 74.5 52.0 5.9 42.2 21.6 9.3 13.2 55.9 84.3 96.1
Llama2-70B 75.2 34.5 5.3 60.2 26.2 7.8 17.5 48.5 79.6 92.2
GPT-4 82.0 52.0 1.5 46.5 15.0 24.0 11.0 50.0 90.5 99.0
NLI Pipeline 58.3 17.6 0.0 82.4 28.9 29.9 12.7 28.4 86.3 98.0
Table 5: Linguistic suitability of generated QA pairs (Angle 2) and hallucinations (hallu). We provide the percent of ratings given by two independent annotators over 100 QA pairs per model. ${}^{\downarrow}$ = lower is better.
6 Results and Discussion
Overall we find that InfoLossQA is a challenging task. While all models have good question-generation/answering abilities, the main difficulty lies in reliably identifying information loss and ensuring the answers are comprehensible for readers. Furthermore, models identify different information loss than humans. The quantitative analysis of the three evaluation angles is given in Tables 4, 5 and 6, accompanied by qualitative examples in Figure 4.
Example 1: Good generation.
$X_{\text{src}}$
: [The aim of this study is] to compare the efficacy of high and low oral doses of haloperidol in the treatment of acute and transient psychotic disorders
$X_{\text{tgt}}$: This study is about testing whether high or low doses of a drug called haloperidol are more effective at treating short-lasting mental illnesses
$Q$: What specific type of mental illness were the patients suffering from? $A$: The patients were suffering from acute and transient psychotic disorders, which are characterized by sudden onset of psychotic symptoms that last for a short period of time.
GPT-4 ✓ Answerable with $X_{\text{src}}$ Vaguely Answerable with $X_{\text{tgt}}$ $E$ is correct $A$ is correct ✓ Givenness $R$ is related ✓ Jargon (+) ✓ Standalone
Example 2: Missing rationale $R$: annotator pointed out that “patients” is an oversimplification of “patients with aBTC,” so it should be highlighted.
$X_{\text{src}}$
: Since the prognosis of advanced biliary tract cancer (aBTC) still remains very poor… Patients with aBTC refractory to at least one regimen of chemotherapies were randomly assigned to receive PPV with low dose CPA … (PPV/CPA, n = 24) or PPV alone (n = 25).
$X_{\text{tgt}}$: Biliary tract cancer (aBTC), a type of liver cancer, has a very poor survival rate. To help improve this, we’re researching new treatments … [W]e gave it to 24 patients and compared them to a group of 25 patients who received only the PPV.
$Q$: Who were the participants in the study?
$A$: The participants were patients with advanced biliary tract cancer (aBTC) who had failed to respond to at least one regimen of chemotherapies.
Mistral ✓ Answerable with $X_{\text{src}}$ Not Answerable with $X_{\text{tgt}}$ $E$ is correct $A$ is correct ✓ Givenness $\sim$ $R$ missing ✓ Jargon (++) ✓ Standalone
Example 3: Even though the QA correctly reveals missing information, the highlighted evidence $E$ does not support the answer.
$X_{\text{src}}$
: This study investigated the effect of combining oral dexamethasone with either nebulized racemic epinephrine or salbutamol compared to bronchodilators alone for the treatment of infants with bronchiolitis. … This was a double-blind, randomized controlled trial on infants (1 to 12 months) who were diagnosed in the emergency department with moderate-to-severe bronchiolitis…
$X_{\text{tgt}}$: This research looked at how well infants with a common lung infection called bronchiolitis responded to different combinations of medicines… The study investigated the effects on babies between 1 to 12 months old…
$Q$: Where were the infants diagnosed with bronchiolitis? $A$: The infants were diagnosed in the emergency department of a hospital.
NLI Pipeline ✓ Answerable with $X_{\text{src}}$ Not Answerable with $X_{\text{tgt}}$ $E$ is wrong $A$ is correct ✓ Givenness $R$ is related ✓ Jargon (++) ✓ Standalone
Example 4: Not an information loss: $A$ does not add any information to $X_{\text{tgt}}$. Givenness: “HbA1c” is not known from the simple text or general knowledge.
$X_{\text{src}}$
: One-hundred participants with a mean±s.d. age of 55.6±10.6 year, body weight of 102.9±18.4 kg and HbA1c of 7.7±1.3% were randomly
$X_{\text{tgt}}$: We had 100 participants on average about 55 years old, weighing about 102.9 kilos, with a blood sugar level of 7.7%. They were placed randomly …
$Q$: What were the ages, weights, and HbA1c levels of the participants in the study?
$A$: The participants had a mean age of 55.6 years, a mean weight of 102.9 kg, and a mean HbA1c level of 7.7%.
Llama-2 ✓ Answerable with $X_{\text{src}}$ Answerable with $X_{\text{tgt}}$ $E$ is correct $A$ is correct ✗ Givenness $\sim$ $R$ missing ✗ Jargon (–) ✓ Standalone
Example 5: Vague question with multiple valid ways of answering; The answer is not standalone because acronym “VS” is only defined in $X_{\text{src}}$.
$X_{\text{src}}$
: … Two experiments were conducted. In Experiment 1, COCs (n = 420) were randomly assigned to four groups: … In Experiment 2, COCs (n = 581) were assigned to the same groups, but those in VS1, VS1 + VS2 and Vitrified groups were sub-divided and exposed to the warming solution for either 1 or 5 min. … In Experiment 2, there was no effect of time in the warming solution…
$X_{\text{tgt}}$: … We did two experiments. In the first experiment, we had 420 [cow egg cells (also called COCs)] separated into four groups: … In the second experiment, we treated 581 COCs the same way but also had some of the COCs from groups 2, 3 and 4 spend either 1 or 5 minutes in a warming solution….
$Q$: What were the details of Experiment 2? $A$: In Experiment 2, 581 COCs were assigned to the same groups as in Experiment 1, but those in VS1, VS1 + VS2 and Vitrified groups were sub-divided and exposed to the warming solution for either 1 or 5 minutes.
Llama-2 $\sim$ Answerable with $X_{\text{src}}$ (Multi.) ✓ Vaguely Answerable with $X_{\text{tgt}}$ $\sim$ $E$ is partial $\sim$ $A$ is partial ✓ Givenness $\sim$ $R$ missing ✗ Jargon (-) ✗ Standalone
Example 6: Both evidence and answer are incomplete because they miss the loading dose of one group.
$X_{\text{src}}$
: … Group L received a loading dose of 0.5 $\mu$g/kg for 10 minutes while group H received 1.0 $\mu$g/kg…
$X_{\text{tgt}}$: … We divided these patients into two groups. One received a lower dosage of the drug, while the other received a higher dosage.
$Q$: What was the specific loading dose used in the study? $A$: The patients in group H received a loading dose of 1.0 $\mu$g/kg.
Mistral $\sim$ Answerable with $X_{\text{src}}$ (Multi.) ✓ Not Answerable with $X_{\text{tgt}}$ $\sim$ $E$ is partial $\sim$ $A$ is partial ✗ Givenness $R$ is related ✗ Jargon (–) ✗ Standalone
Figure 4: Qualitative examples demonstrating each error case (excluding hallucinations for space constraints).
Models generate valid questions but not all are about an information loss (Table 4).
We find that the majority of questions is answerable with $X_{\text{src}}$, showing that all models are good at document-grounded question generation. However, a sizable number of questions is not scoped to any particular information loss, having multiple valid answers (7.8%–19.6%). For example, “What are the details of the experiments?” (Ex. 5 in Figure 4). Furthermore, a small number of questions is unanswerable with $X_{\text{src}}$. Qualitatively, these questions tend to ask for background explanations of technical concepts which are not about an information loss.
Despite most questions being valid, all models could improve at generating questions that are not or only vaguely answerable with $X_{\text{tgt}}$. See Ex. 4 (Figure 4) for an example where the generated QA does not add any new information. Compared to open-source LLMs, the fact-based NLI pipeline is substantially better at identifying unanswerable questions demonstrating this approach is a promising direction for identifying information loss.
Considering answer accuracy, GPT-4 provides near perfect results both in extracting the correct evidence span (98% correct) and in rewriting it into a full answer (94% correct). Surprisingly, Mistral and Llama2 more often generate correct abstractive answers than extracting the right evidence. As the abstractive answer should follow from the evidence, those LLM responses are not internally consistent. Lastly, across all models, we only observe a small rate of hallucinations (worst: Llama2, 7.8%).
Models often fail to localize QAs in the simple text (Table 5).
In terms of Rationale Localization, too often models do not produce a rationale span in $X_{\text{tgt}}$, which shows a general limitation of models to correctly discern between deletion and oversimplification. For reference: 25% of human QAs are deletions, whereas the models classify more than 50% of QAs as deletions (Table A.2). Annotators have substantial agreement for when there should be a rationale span (Table 3), so this could be alleviated by model fine-tuning.
QAs are well-framed, with room for improvement in terms of jargon (Table 5).
With respect to the Givenness constraint, the end-to-end models provide well-framed questions that only contain concepts familiar to readers and do not have any answer leakage. This result is on par or slightly better than recent QUD parsers Wu et al. (2023).
Concerning the answers, most can be understood without referring to the original text, i.e., they are Standalone. Qualitatively, undefined abbreviations are frequent reasons for answers to not be standalone (Figure 4) which could be solved with a dedicated decontextualization step Choi et al. (2021).
Lastly, a large portion of answers contains unexplained Jargon (–, best NLI 28.4%, worst Mistral 55.9%). While all models were instructed to generate simple answers, we assume that the task has too many constraints to be expressed in an end-to-end prompt. By decoupling the identification of information loss and QA generation, the NLI pipeline can better follow this constraint.
How are errors distributed across document parts?
Given that some RCT sections are more technical than others, it is conceivable that models do better in some sections than in others. We plot the percentage of good responses per model, section and evaluation criterion in Figure A.1. Contrary to our expectation, there are no notable differences across sections, with only a slight trend for reduced Givenness and Jargon for QAs localized in the results and conclusion section. Qualitatively, this is often due to questions asking about the statistical significance of the results, which annotators deemed to be an unfamiliar concept for lay readers.
Model Full (normal-$\uparrow$) Partial No (normal-$\downarrow$)
Mistral-7B 14.6 4.0 81.4
Llama2-70B 16.1 4.3 79.6
GPT-4 28.7 4.0 67.3
NLI Pipeline 22.6 21.5 56.0
Table 6: Recall (%) of human-written QA pairs across 55 documents (554 QA pairs). A human QA pair can either be fully, partially or not recalled.
Models have a low recall of human QAs (Table 6).
The model with highest recall is the NLI pipeline (44% of human QA at least partially recalled). However, this model also generated significantly more questions per document than humans (16.5 vs. 9.6, Table A.2). Qualitatively, the NLI pipeline generates questions which each cover a smaller part of missing content, whereas humans aggregate related missing information into one larger question (see Figure A.2 for an example). This is also corroborated by the NLI pipeline having the highest partial recall of human QAs. Exciting directions for future work will be to (1) get models to apply similar standards as humans at what constitutes an information loss, and (2), to study how to aggregate/summarize smaller-scope questions into larger questions.
7 Related Work
Deletion in text simplification.
Deletion is a strategy professional editors commonly take to render text more accessible to lay readers Petersen and Ostendorf (2007); Xu et al. (2015); Zhong et al. (2020); Yamaguchi et al. (2023). Of note are several studies that analyzed deletions in simplification corpora and model outputs. Devaraj et al. (2022) devised a typology of deletion errors, distinguishing between meaning-preserving and meaning-altering deletion. They found that, at the sentence level, even professional simplifications include overdeletion errors and that models are prone to learn these from the training data. Agrawal and Carpuat (2023) conducted user studies to analyze the effect of simplification on document-level reading comprehension. They found that deletions are a major factor for diminished reading comprehension. Overall, these studies highlight the importance of detecting and mitigating deletions. We directly build on this line of work by providing the first annotated dataset of information loss and a QA-based framework to identify and resurface this information to lay readers. Contrary to above work, we do not restrict information loss to deletions, but take a broader perspective where critical information can also be lost through oversimplification.
Question generation (QG).
QG is the task of generating a question given a textual input Du et al. (2017); Zhou et al. (2018). This line of work has seen many successful applications, e.g., in improving QA Rajpurkar et al. (2016); Duan et al. (2017). While early work considered answer-aware QG with factoid answers, the focus shifted to natural, information-seeking questions Kwiatkowski et al. (2019); Dasigi et al. (2021).
Recent work investigates the generation of curiosity-driven and inquisitive questions, which are questions asked by humans with the goal to enrich their knowledge Ko et al. (2020); Scialom and Staiano (2020); Meng et al. (2023). A particular kind of question in this domain are clarification questions Rao and Daumé III (2018); Majumder et al. (2021). We consider questions with a similar pragmatic goal as clarification questions: namely to identify information which is missing or vague given a textual context. Prior work has demonstrated the merits of QG to represent missing information for tasks like decontextualization Newman et al. (2023). But we draw attention to a particularity of QG in text simplification is that lay people reading a simplified text have unknown unknowns, preventing them from asking these clarification questions. Compared to experts which have expectations of what is included in a technical text, lay readers have difficulty identifying this missing information because they lack the necessary “disciplinary knowledge” August et al. (2023). From a technical perspective, we are concerned with the subset of clarification questions that are fully answerable with the original text, but not/vaguely answerable with the simplified text.
Our work is closely related to DiffQG Cole et al. (2023). DiffQG aims to generate questions that describe factual changes between two revisions of a Wikipedia passage. Compared to DiffQG, we consider simplification-induced changes where answers to questions are not necessarily different but vague. Furthermore, our task requires document-level comparisons, answers that can range from simple factoid to complex multi-sentence answers, and tailoring the QA to laypeople.
QA for evaluation and entailment.
Closely related is the idea of using question-answering to evaluate summarization and simplification Mani et al. (2002); Agrawal and Carpuat (2023). The advances in automatic QG (see above) gave rise to the cross-questioning paradigm to assess factuality and relevance of generations which has been extensively studied (Wang et al., 2020; Durmus et al., 2020; Deutsch et al., 2021, inter alia). This line of work differs from InfoLossQA in two important aspects. First, there is a strong focus on noun-phrase and entity-centric QA. Our proposed QAs invite various types of answers. In addition, the QAs are meant for consumption by end users and therefore must satisfy additional linguistic criteria (Section 2). Second, while it would be conceivable to use the cross-questioning paradigm to obtain candidate QAs, Kamoi et al. (2023) identified error-propagation in the QG stage as a fundamental limitation that make QA-level answerability judgments unreliable. Kamoi et al. (2023) therefore advocate for NLI-based approaches Laban et al. (2022); Rodriguez et al. (2023), which our pipeline-system draws inspiration from.
8 Conclusion
To mitigate overdeletion and vagueness in text simplification, we propose InfoLossQA: A task and dataset that aims to describe and recover simplification-induced information loss as reader-focused QA pairs. Our analyses show that QAs following the QUD theory provide a rich description of what information is lost. We also establish automatic baselines for the task and propose a rigorous evaluation framework considering correctness of QA pairs and user-centric constraints. We find that today’s methods struggle to reliably identify information loss and to apply similar standard as humans as to what constitutes information loss. Looking ahead, we believe that it is exciting to explore the presented methods in interactive simplification tools Fok et al. (2023) and for quality assessment of automatic simplifications.
Limitations
Our motivation and vision for InfoLossQA is to help users to deepen their understanding of a text. However, we focus on the technical dimensions of this goal: establishing a dataset, an evaluation framework, and developing and evaluating baselines. Testing the effects of the proposed QA on end-user comprehension is an important direction for future work in interactive text simplification.
Furthermore, while we believe that the proposed approach is generalizable, our experiments are confined to one language (English), text genre (abstracts of medical publications) and simplification style (GPT-4 simplification). Future work could assess the versatility of this framework under different conditions. For instance, we considered simplifications at one level of compression, but readers may prefer different simplification lengths and language levels Xu et al. (2015). As we alter the length of simplifications, it becomes important to understand (a) how effective models are at identifying information loss, and (b) how the distribution of information loss and associated QAs changes.
Lastly, our proposed evaluation framework currently relies on human annotators to judge the quality of model outputs. Any kind of human evaluation comes at a significant cost which may limit further studies on better modeling. Therefore, future work could develop automatic metrics for the task. We release the human judgments for model outputs to facilitate this direction.
Acknowledgments
We thank Keziah Kaylyn Reina, Kathryn Kazanas and Karim Villaescusa F. for their annotation and evaluation effort, David Heinemann for help with the annotation interface, and Ritvik Renikunta for help with the initial design of the end-to-end prompts. We also thank Eunsol Choi for feedback on this paper, and Juan Diego Rodriguez and Manya Wadhwa for useful discussions. This research is partially supported by NSF CAREER Award IIS-2145479 and Good Systems, [^14] https://goodsystems.utexas.edu a UT Austin Grand Challenge to develop responsible AI technologies. Trienes was supported by the Cancer Research Center Cologne Essen (CCCE), the Federal Ministry of Education and Research (BMBF) and by a fellowship within the IFI programme of the German Academic Exchange Service (DAAD). Wallace was supported in this work by the National Institutes of Health (NIH), grant R01LM012086, and by the National Science Foundation (NSF), grant 1750978. Xu is supported in part by NSF awards IIS-2144493 and IIS-2112633.
Appendix A Appendix
Description Prompt Decoding Parameters
Simplification. Simplifying technical RCT abstracts ($X_{\text{src}}$) into a plain language version ($X_{\text{tgt}}$).
Figure A.3 model=gpt-4-0613
temperature=1.0
max_tokens=1024
top_p=1
frequency_penalty=0
presence_penalty=0
Few-shot question classifier. Categorizing questions according to the typology of Cao and Wang (2021). This classifier obtains an F1 score of 0.87 on a manually labeled validation set of 50 questions in InfoLossQA.
Figure A.4 model=gpt-4-0613
temperature=0
max_tokens=1024
top_p=1
frequency_penalty=0
presence_penalty=0
NLI pipeline: fact extraction. Extracting atomic facts from $X_{\text{src}}$. These facts are classified for entailment with $X_{\text{tgt}}$ with a standard NLI model Laurer et al. (2024).
Figure A.5 model=gpt-4-0613
temperature=0
max_tokens=512
top_p=1
frequency_penalty=0
presence_penalty=0
NLI pipeline: QA-generation. Generating QA pairs based on $X_{\text{src}}$, $X_{\text{tgt}}$ and the list of facts with neutral outcome of the NLI classifier.
Figure A.6 model=gpt-4-0613
temperature=1
max_tokens=4096
top_p=1
frequency_penalty=0
presence_penalty=0
GPT-4. End-to-end prompt to detect information loss and generate QA pairs.
Figure A.7 model=gpt-4-0613
temperature=0
max_tokens=2048
top_p=1
frequency_penalty=0
presence_penalty=0
Mistral-7B. End-to-end prompt to detect information loss and generate QA pairs. Inference with the Huggingface transformers library.
Figure A.8 model=Mistral-7B-Instruct-v0.1
do_sample=False
temperature=0
max_tokens=2048
top_p=1
top_k=1
repetition_penalty=1
Llama2-70B. End-to-end prompt to detect information loss and generate QA pairs. Inference on API of Together.AI.
Figure A.9 model=llama-2-70b-chat
max_tokens=None
temperature=0
top_p=1
top_k=1
repetition_penalty=1
Table A.1: LLM prompts and decoding parameters.
Annotator / Model Docs. QA % Over % Del QA/doc $|Q|$ $|A|$ $|E|$ $|R|$
Annotator 1 48 192 90.1 9.9 4.1 10.8 23.8 12.2 7.6
Annotator 2 75 425 71.8 28.2 5.7 12.2 33.4 15.5 9.0
Annotator 3 85 383 68.7 31.3 4.7 10.9 30.2 14.4 8.6
All (micro avg.) 104 1000 74.1 25.9 9.6 11.4 30.3 14.4 8.5
Mistral-7B 104 507 45.2 54.8 4.9 11.6 26.2 22.3 18.1
Llama2-70B 104 681 38.2 61.8 6.7 12.9 30.6 20.4 15.7
GPT-4 104 477 48.4 51.6 4.6 14.4 33.7 25.4 20.6
NLI (GPT-4) 104 1699 100.0 16.5 14.5 24.6 34.7
Table A.2: Summary statistics of human-written QAs and model predictions. Over = Oversimplification; Del = Deletion; Length of question $Q$, answer $A$, evidence spans $E\in X_{\text{src}}$ and rationale spans $R\in X_{\text{tgt}}$ is given in tokens.
Error Mistral-7B Llama2-70B GPT-4 NLI Pipeline
Deletion (Total) 278 421 246 1699
% Valid (no error) 61.9 87.2 100 100
% Spurious $R$ 38.1 1.0 - -
% Invalid $E$ 11.9 - -
Oversimplification (Total) 229 260 231 0
% Valid (no error) 67.2 78.8 100 -
% Invalid $R$ 28.8 5.8 - -
% Missing $R$ 2.2 6.2 - -
% Invalid $E$ 1.7 2.3 - -
% Invalid $E$ + Invalid $R$ 5.8 - -
% Invalid $E$ + Missing $R$ 1.2 - -
Table A.3: Analysis of LLM generation errors grouped by deletion and oversimplification. For each category, the total number of QA pairs is given with the percent of QAs per error. Spurious $R$: QAs classified as deletions should not have a rationale span. Invalid $E$/$R$: the span cited by the model is not a valid substring of $X_{\text{src}}$/$X_{\text{tgt}}$.
QA Category % Definition and Examples
Motivation 5.0%
Why was the study conducted?
Q: What was the motivation for this study to investigate the dosage of haloperidol to address nausea and vomiting after surgery?
A: There is evidence that a small dose of haloperidol can help prevent…
Condition 5.8% What condition does this paper study?
– Inclusion Criteria 3.3%
What were the specific inclusion criteria for participants?
Q: What criteria did the researchers use to select eligible participants?
A: The researchers gathered participants between the ages of 3-18, …
– Other 2.5%
Other condition-related question.
Q: What kind of hip surgery did the participants receive?
A: The participants were getting hip surgery under the subarachnoid block…
Goal 10.8%
What did the paper want to find out?
Q: What about the nicotine vaccine could potentially help smokers quit?
A: The nicotine vaccine boosts antibody concentrations, which helps to stimulate…
Usual Treatment 1.7%
How is the condition usually treated?
Q: What type of active deep brain stimulation was used in previous trials?
A: Previous trials found a positive impact of unilateral (only applied to one side of the brain) active deep brain stimulation on symptoms of Parkinson’s disease.
New Treatment 4.2%
What were the new treatment(s), if any this paper looked into?
Q: What type of formoterol is being observed in the study?
A: Long-acting beta2-agonist formoterol, which is a drug used to treat asthma…
Method 53.3% What did the paper do?
– Outcome (Tool) 14.2%
What tools/procedures were used to measure the effects of interventions?
Q: How was sleep quality and life quality measured for participants?
A: Sleep quality and life quality were assessed using the Sleep-Apnoea-Quality-of-Life-Index…
– Study Protocol 10.8%
What was the protocol of the study?
Q: What was the setting and design of the study?
A: The study was an open (both participants and researchers knew who was assigned which drops), cross-over (the participants receive both treatments …), comparative study (comparing the two drops).
– Quantity 7.5%
With what dosage/quantity/frequency were the interventions performed?
Q: How much lurasidone was given to the patients in the trial?
A: Patients were given 40 to 80 milligrams of lurasidone per day, given flexibly.
– Population 7.5%
What were the demographics of the patients in the study?
Q: What was the age range of the patients?
A: 52-85 years.
– Analysis Technique 5.8%
How were the results analyzed?
Q: How were the results analyzed?
A: The results of the test were analyzed using a t-test and…
– Other 7.5%
Other method-related question.
Q: How often were participants checked during the study?
A: Participants were assessed on their first visit, on the 7th day or their first follow-up, and on the 14th day or their second follow-up.
Results 18.3%
What did the paper find?
Q: How did the number of evens in the Tamoxifen group compare to the other group?
A: The Tamoxifen group had a total of 39 events, while the other…
Contrasting Results 0.8%
Are the findings different depending on a person’s demographics?
Q: How did selenium levels of pregnant women in the UK vary from those of pregnant women in other countries?
A: The selenium (mineral) levels of pregnant women in the UK were lower compared to those of pregnant women in other countries…
Table A.4: Classification of 120 randomly sampled human QA pairs in InfoLossQA according to the pragmatic goals of questions in the Paper Plain Key Question Index developed for lay readers of RCTs August et al. (2023). Frequency of questions is given in percent (%). (Sub-)Categories marked with were added by us to give a more fine-grained view on the questions types.
Figure A.1: Percentage of QAs with a good rating grouped by model, evaluation criterion and RCT section (BG =Background; ME = Methods; RE = Results; CO = Conclusion; n/a for RCTs without sections). Criteria with multiple response options were binarized to “good” vs. “not good.”
Figure A.2: Full example comparing QAs generated by the NLI pipeline, humans and GPT-4. Colors indicate semantic alignment between questions. We observe different tendencies for question compositionality: while humans aggregate several missing information into higher-level questions, the NLI pipeline has a tendency to generate more restricted questions, even though in aggregate they cover a similar content.
Please simplify the following technical abstract into plain language that an average adult would understand. If the abstract has sections, keep them.
{{ abstract }}
Figure A.3: GPT-4 prompt to simplify the RCT abstracts. Colors indicate chat roles: system and user.
You are a helpful assistant to classify text into categories. ## Instructions You are asked to classify questions according to an ontology of question types. The question type reflects the nature of the question. It is NOT determined by the interrogative word of the question. There are 10 question types in total. The definition for each type is shown below. Please select the question type which is most likely for a given question. Only output the category title, not the description. ## Question Types 1. VERIFICATION: Asking for the truthfulness of an event or a concept. - Was the study double-blinded? - Was there a trend towards smaller increases in macular pigment for subjects with high baseline values? 2. DISJUNCTION: Asking for the true one given multiple events or concepts, where comparison among options is not needed. - no example available, match by the description 3. CONCEPT: Asking for a definition of an event or a concept. - What kind of result is being measured in this study? - What were the main inclusion criteria for this study? - What does the WCJ-III test specifically measure? - Which areas of the brain were analyzed? 4. EXTENT: Asking for the extent or quantity of an event or a concept. - How reliable are these results? - How long were the participants observed? - How much ibuprofen was in the small dose? - How many young and old people participated in the study? 5. EXAMPLE: Asking for example(s) or instance(s) of an event or a concept. - What kind of conditions cause corneal neovascularization? 6. COMPARISON: Asking for comparison among multiple events or concepts. - On what results did the control group do better than the intervention group? - How did headache of participants receiving ibuprofen compare to those participants that received a placebo? - How more effective was the arm cranking exercise with and without electrical muscle stimulation? 7. CAUSE: Asking for the cause or reason for an event or a concept. - What motivates this study? - Why is EMS being investigated? 8. CONSEQUENCE: Asking for the consequences or results of an event. - What was the effect of using ibuprofen to treat headaches? - What were the main findings of the study? - What does this study tell us about arm cranking with electrical muscle stimulation? 9. PROCEDURAL: Asking for the procedures, tools, or methods by which a certain outcome is achieved. - What questionnaire was used for this study? - What kind of lab tests were done? - How were the patients assigned to a group? - How were the different drugs administered to participants? 10. JUDGMENTAL: Asking for the opinions of the answerers own. - no example available, match by the description
## Instances to classify {for i, question in enumerate(questions) %}     {{ i+1 }}. {{ question }} {endfor %}
Figure A.4: Few-shot prompt for GPT-4 to classify questions according to the taxonomy of Cao and Wang (2021). Colors indicate chat roles: system and user.
Statement: {{ sent }} List all the facts we explicitly know from the statement. Make each fact as atomic as possible.
Figure A.5: GPT-4 prompt to decompose a sentence into a list of atomic facts.
## Original {{ original }} ## Simplification {{ simplification }} ## Missing facts {{ facts }} The above facts are missing from the simplified text ("Simplification"). For each fact, please write a question-answer pair that would elicit the missing information from the original text ("Original"). Phrase the question in such a way that a reader can understand it without having seen the original text. It should only contain concepts (entities, events, or states) that were mentioned in the simple text, or concepts that have not been directly mentioned but are generally known or inferable from mentioned ones. The answer should be understandable by an average adult, so please explain technical jargon if necessary. Make each question-answer pair as specific as possible and make sure that they are independent of each other. Ask only about one information unit at a time. Do this for all facts, and format your output as follows: - Fact: - Question: - Answer:
Figure A.6: GPT-4 prompt to generate an InfoLossQA pair given original, simplification and a missing fact.
You are an expert annotator for outputs of text simplification systems. This annotation task is to identify pieces of information that were lost in the simplification process. You will be given two texts: the original and a simplification. Assume that a lay reader only sees the simplification. Identify all things which the reader can NOT learn from the simplification but that they could have learnt from the original. Please classify each unit of information into one of the following two categories: 1. Omissions: Pieces of information which were not included in the simplification. 2. Imprecise/fuzzy concepts: Pieces of information which are included in the simplification, but that have been simplified to the extent that they became imprecise or completely lost their original meaning. Afterwards, please write a question-answer pair that would elicit the missing information from the original text. Phrase the question in such a way that a reader can understand it without having seen the original text. It should only contain concepts (entities, events, or states) that were mentioned in the simple text, or concepts that have not been directly mentioned but are generally known or inferable from mentioned ones. The answer should be understandable by an average adult, so please explain technical jargon if necessary. Make each question-answer pair as specific as possible and make sure that they are independent of each other. Ask only about one information unit at a time. A useful heuristic to decide between omissions and imprecise/fuzzy concepts is to see if the question-answer pair clarifies/expands some topic which is discussed in the simplification. If so, classify it as imprecise/fuzzy concepts, otherwise omission. Adhere to this output format: - Original: "<span in original text>" - Rationale: <a short rationale that describes what makes this an information loss> - Question: <the question that elicits missing information> - Answer: <the answer that provides the missing information from the original span> For fuzzy/imprecise concepts, please also indicate the corresponding span in the simplified text. Here is an example. ## Original This study evaluates the effects of vitamin D3 on disease activity and quality of life in ulcerative colitis (UC) patients with hypovitaminosis D. The study was a prospective double-blinded, randomized trial conducted at Community Regional Medical Center, Fresno, CA from 2012-2013. Patients with UC and a serum 25(OH)D level <30 ng/ml were eligible for the study. Enrolled subjects were randomized to receive either 2,000 lU or 4,000 IU of oral vitamin D3 daily for a total of 90 days. The Short IBD Questionnaire (SIBDQ) for quality of life, the Partial Mayo Score for UC disease activity and serum lab tests were compared between the two treatment groups. ## Simplification This study looks at whether taking vitamin D3 can affect a particular form of bowel disease (ulcerative colitis) and improve the lives of patients with low levels of vitamin D. The study was carried out at a medical center in Fresno, California, between 2012 and 2013. Patients with this disease and low vitamin D levels were included. Participants were randomly given either 2,000 IU or 4,000 IU of oral vitamin D3 daily for 90 days. Researchers used a survey about participants quality of life and conducted lab tests between the two groups. ## Omissions - Original: "Partial Mayo Score for UC disease activity" - Rationale: the simplification does not mention that the Partial Mayo Score was used. - Question: Besides quality of life survey and the serum tests, what other measures did the researchers use? - Answer: The researchers also tracked the partial mayo for UC diseases. UC stands for ulcerative colitis and is an inflammatory bowel disease. The mayo score indicates how active or severe the disease is. ## Imprecise/fuzzy concepts - Original: "The study was a prospective double-blinded, randomized trial" - Simplification: "The study" - Rationale: The simplification does not explain the design of the study, it only mentions that it was a "study". - Question: How did the study control for bias? - Answer: The study was double-blinded, so that neither the researcher nor the participants knew which treatment each participant received, and it was randomized, meaning the participants were randomly assigned one of the treatments. - Original: "a serum 25(OH)D level <30 ng/ml" - Simplification: "low vitamin D levels" - Rationale: the simplification does not explain the inclusion criteria, namely how low the vitamin D levels of eligible patients were. - Question: How low were the vitamin D levels in patients that were included in the study? - Answer: Participants in the study all had less than 30ng/ml of vitamin D, which is below the minimum recommendation of vitamin D levels in the body. - Original: "Short IBD Questionnaire (SIBDQ) for quality of life" - Simplification: "a survey about participants quality of life" - Rationale: the specific type of survey (SIBDQ) is not mentioned in the simplification - Question: What survey was used to measure participants quality of life? - Answer: The Short Inflammatory Bowel Disease Questionnaire (SIBDQ) was used which gives insights about the physical, social, and emotional status of patients with bowel diseases.
## Original {{ original }} ## Simplification {{ simplification }}
Figure A.7: End-to-end system prompt for GPT-4. Colors indicate chat roles: system and user.
You are an expert annotator for outputs of text simplification systems. This annotation task is to identify pieces of information that were lost in the simplification process. You will be given two texts: the original and a simplification. Assume that a lay reader only sees the simplification. Identify all things which the reader can NOT learn from the simplification but that they could have learnt from the original. Please classify each unit of information into one of the following two categories: 1. Omissions: Pieces of information which were not included in the simplification. 2. Imprecise/fuzzy concepts: Pieces of information which are included in the simplification, but that have been simplified to the extent that they became imprecise or completely lost their original meaning. Afterwards, please write a question-answer pair that would elicit the missing information from the original text. Phrase the question in such a way that a reader can understand it without having seen the original text. It should only contain concepts (entities, events, or states) that were mentioned in the simple text, or concepts that have not been directly mentioned but are generally known or inferable from mentioned ones. The answer should be understandable by an average adult, so please explain technical jargon if necessary. Make each question-answer pair as specific as possible. Ask only about one information unit at a time. A useful heuristic to decide between omissions and imprecise/fuzzy concepts is to see if the question-answer pair clarifies/expands some topic which is discussed in the simplification. If so, classify it as imprecise/fuzzy concepts, otherwise omission. Adhere to this output format: - Original Fragment: <exact span in original> - Simplification Fragment: null for Omissions OR <exact span in simplification> for Imprecise/fuzzy concepts - Rationale: <a short rationale that describes what makes this an information loss> - Question: <the question that elicits missing information> - Answer: <the answer that provides the missing information from the original span> Strictly include the above elements. There can be many omissions and imprecise concepts, so be exhaustive. Use simple language in the answer. Cite the original/simplification span EXACTLY so that span == text[text.index(span):text.index(span)+len(span)]. ## Original This study evaluates the effects of vitamin D3 on disease activity and quality of life in ulcerative colitis (UC) patients with hypovitaminosis D. The study was a prospective double-blinded, randomized trial conducted at Community Regional Medical Center, Fresno, CA from 2012-2013. Patients with UC and a serum 25(OH)D level <30 ng/ml were eligible for the study. Enrolled subjects were randomized to receive either 2,000 lU or 4,000 IU of oral vitamin D3 daily for a total of 90 days. The Short IBD Questionnaire (SIBDQ) for quality of life, the Partial Mayo Score for UC disease activity and serum lab tests were compared between the two treatment groups. ## Simplification This study looks at whether taking vitamin D3 can affect a particular form of bowel disease (ulcerative colitis) and improve the lives of patients with low levels of vitamin D. The study was carried out at a medical center in Fresno, California, between 2012 and 2013. Patients with this disease and low vitamin D levels were included. Participants were randomly given either 2,000 IU or 4,000 IU of oral vitamin D3 daily for 90 days. Researchers used a survey about participants quality of life and conducted lab tests between the two groups.
## Omissions - Original Fragment: "Partial Mayo Score for UC disease activity" - Simplification Fragment: null - Rationale: the simplification does not mention that the Partial Mayo Score was used. - Question: Besides quality of life survey and the serum tests, what other measures did the researchers use? - Answer: The researchers also tracked the partial mayo for UC diseases. UC stands for ulcerative colitis and is an inflammatory bowel disease. The mayo score indicates how active or severe the disease is. ## Imprecise/fuzzy concepts - Original Fragment: "The study was a prospective double-blinded, randomized trial" - Simplification Fragment: "The study" - Rationale: The simplification does not explain the design of the study, it only mentions that it was a "study". - Question: How did the study control for bias? - Answer: The study was double-blinded, so that neither the researcher nor the participants knew which treatment each participant received, and it was randomized, meaning the participants were randomly assigned one of the treatments. - Original Fragment: "a serum 25(OH)D level <30 ng/ml" - Simplification Fragment: "low vitamin D levels" - Rationale: the simplification does not explain the inclusion criteria, namely how low the vitamin D levels of eligible patients were. - Question: How low were the vitamin D levels in patients that were included in the study? - Answer: Participants in the study all had less than 30ng/ml of vitamin D, which is below the minimum recommendation of vitamin D levels in the body. - Original Fragment: "Short IBD Questionnaire (SIBDQ) for quality of life" - Simplification Fragment: "a survey about participants quality of life" - Rationale: the specific type of survey (SIBDQ) is not mentioned in the simplification - Question: What survey was used to measure participants quality of life? - Answer: The Short Inflammatory Bowel Disease Questionnaire (SIBDQ) was used which gives insights about the physical, social, and emotional status of patients with bowel diseases.
## Original {{ original }} ## Simplification {{ simplification }}
Figure A.8: End-to-end system prompt for Mistral. Colors indicate chat roles: user and assistant.
You are an expert annotator for outputs of text simplification systems. You are given an example, a sample to annotate and detailed instructions.
### Original This study evaluates the effects of vitamin D3 on disease activity and quality of life in ulcerative colitis (UC) patients with hypovitaminosis D. The study was a prospective double-blinded, randomized trial conducted at Community Regional Medical Center, Fresno, CA from 2012-2013. Patients with UC and a serum 25(OH)D level <30 ng/ml were eligible for the study. Enrolled subjects were randomized to receive either 2,000 lU or 4,000 IU of oral vitamin D3 daily for a total of 90 days. The Short IBD Questionnaire (SIBDQ) for quality of life, the Partial Mayo Score for UC disease activity and serum lab tests were compared between the two treatment groups. ### Simplification This study looks at whether taking vitamin D3 can affect a particular form of bowel disease (ulcerative colitis) and improve the lives of patients with low levels of vitamin D. The study was carried out at a medical center in Fresno, California, between 2012 and 2013. Patients with this disease and low vitamin D levels were included. Participants were randomly given either 2,000 IU or 4,000 IU of oral vitamin D3 daily for 90 days. Researchers used a survey about participants quality of life and conducted lab tests between the two groups. ### Missing information in simplification
[     {         "category": "imprecise/fuzzy concepts",         "original": "The study was a prospective double-blinded, randomized trial",         "simplification": "The study",         "rationale": "The simplification does not explain the design of the study, it only mentions that it was a study’.",         "question": "How did the study control for bias?",         "answer": "The study was double-blinded, so that neither the researcher nor the participants knew which treatment each participant received, and it was randomized, meaning the participants were randomly assigned one of the treatments.",     },     {         "category": "imprecise/fuzzy concepts",         "original": "a serum 25(OH)D level <30 ng/ml",         "simplification": "low vitamin D levels",         "rationale": "The simplification does not explain the inclusion criteria, namely how low the vitamin D levels of eligible patients were.",         "question": "How low were the vitamin D levels in patients that were included in the study?",         "answer": "Participants in the study all had less than 30ng/ml of vitamin D, which is below the minimum recommendation of vitamin D levels in the body."     },     {         "category": "imprecise/fuzzy concepts",         "original": "Short IBD Questionnaire (SIBDQ) for quality of life",         "simplification": "a survey about participants quality of life",         "rationale": "The specific type of survey (SIBDQ) is not mentioned in the simplification",         "question": "What survey was used to measure participants quality of life?",         "answer": "The Short Inflammatory Bowel Disease Questionnaire (SIBDQ) was used which gives insights about the physical, social, and emotional status of patients with bowel diseases.",     },     {         "category": "omission",         "original": "Partial Mayo Score for UC disease activity",         "simplification": null,         "rationale": "The simplification does not mention that the Partial Mayo Score was used.",         "question": "Besides quality of life survey and the serum tests, what other measures did the researchers use?",         "answer": "The researchers also tracked the partial mayo for UC diseases. UC stands for ulcerative colitis and is an inflammatory bowel disease. The mayo score indicates how active or severe the disease is.",     }, ]
### Original {{ original }} ### Simplification {{ simplification }} ### Instructions This annotation task is to identify pieces of information that were lost in the simplification process. Annotate the sample above. Assume that a lay reader only sees the simplification. Identify all things which the reader can NOT learn from the simplification but that they could have learnt from the original. Please classify each unit of information into one of the following two categories: 1. Omissions: Pieces of information which were not included in the simplification. 2. Imprecise/fuzzy concepts: Pieces of information which are included in the simplification, but that have been simplified to the extent that they became imprecise or completely lost their original meaning. Afterwards, please write a question-answer pair that would elicit the missing information from the original text. Phrase the question in such a way that a reader can understand it without having seen the original text. It should only contain concepts (entities, events, or states) that were mentioned in the simple text, or concepts that have not been directly mentioned but are generally known or inferable from mentioned ones. The answer should be understandable by an average adult, so please explain technical jargon if necessary. Make each question-answer pair as specific as possible and make sure that they are independent of each other. Ask only about one information unit at a time. A useful heuristic to decide between omissions and imprecise/fuzzy concepts is to see if the question-answer pair clarifies/expands some topic which is discussed in the simplification. If so, classify it as imprecise/fuzzy concepts, otherwise omission. Include following elements in your annotation: - Original: <exact span in original> - Simplification: <exact span in simplification> or null for omissions - Rationale: <a short rationale that describes what makes this an information loss> - Question: <the question that elicits missing information> - Answer: <the answer that provides the missing information from the original span> Strictly follow the above json format. There can be many omissions and imprecise concepts, so be exhaustive. Use simple language in the answer. Cite the original/simplification span EXACTLY so that span == text[text.index(span):text.index(span)+len(span)]. Output ONLY the json! ### Missing information in simplification
Figure A.9: End-to-end system prompt for Llama2. Colors indicate chat roles: system, user and assistant.
Appendix B Annotation Guidelines
Introduction.
Text simplification aims to rewrite a complex text into a simpler version that can be understood by a lay audience. When simplifying, editors decide what and how to simplify, often omitting content which is deemed less important or too technical. However, this can deny readers access to potentially useful information and the opportunity to learn new terms and concepts.
Goal.
We aim to identify instances where information is lost and to recover it through Question-Answer (QA) pairs. We hypothesize that readers can better understand the text by referring to the simplified version and the associated QA pairs.
The data.
We work with abstracts of randomized controlled trials (RCTs) along with their automatically generated simplifications. RCTs are scientific experiments testing the efficacy of clinical interventions like new drugs, treatments, or diagnostic methods. They typically involve recruiting patients and dividing them into an experimental group (receiving the intervention) and a control group (not receiving it).
Annotation task.
Your task involves the following steps:
  1. 1.
    Read both the original and simplified text
  2. 2.
    Compare the two texts and highlight
    1. (a)
      Deletions from the original, and
    2. (b)
      Oversimplifications in the simplification
  3. 3.
    For each highlight, write a QA pair that re-introduces the information in lay language.
You can find two examples below.
Deletions.
Pieces of information which were not included in the simplification. The questions should allow readers to reveal the omitted content.
Original: First, nine healthy young men performed two ➀ 20-min arm-cranking trials ➁ at 50% VO2-max with and without EMS applied to the lower limbs.
Simplified: First, they had nine healthy young men do the arm-cranking exercise with and without the EMS added to the lower body.
Instance 1:
Q: For how long is the exercise applied?
A: Participants do the exercise twice for 20 minutes.
Instance 2:
Q: At what intensity is the exercise applied?
A: The exercise is applied at 50% VO2-max.
Oversimplifications.
Pieces of information that were simplified to the extent that they became too imprecise or completely lost their original meaning. The questions should clarify those concepts and restore their original meaning. For this category we highlight both the simplification and the original text that elicited the simplification.
Original: The flow-mediated vasodilation (FMD) of the right brachial artery was measured using a high-resolution ultrasound device.
Simplified: They used a high-resolution ultrasound to measure how well the main artery in the arm was working.
Q: How do researchers measure how well the intervention (EMS) works?
A: The researchers measure to what extent the main artery of the arm widens, which is called flow-mediated vasodilation (FMD).
Scenario grounding: deciding what is important enough.
  • Do role play. Imagine you are someone who could benefit from understanding the RCT. For example, a patient who has a condition that this RCT addresses.
  • When in doubt, try to take an inclusive perspective. People can always decide a question is not relevant to them.
  • For numerical results like p-values or Z-scores, make a judgment if they are necessary for a correct understanding of the RCT. Do not interpret these results, rather rely on the authors interpretation of the values.
Tips and other notes.
  • When is something a deletion vs. oversimplification? A useful heuristic is to see if you could “attach” the omitted information somewhere to the simplification. If so, it is likely an oversimplification.
  • You are free to use a search engine for writing the answers. Please only use high quality sources.
  • The text may refer to the same concept multiple times. Please highlight all occurrences, and copy-paste the QA. Add any comments, questions or concerns in the comment field.
  • Ignore other artifacts in the simplifications that are not about an information loss, including factual errors, under-simplifications (i.e., when something is still too complicated) and unnecessary information (e.g., the registration number of a trial).
Checklist for a good QA pair.
  • Address an information gap between the original and the simplification.
  • A question should be self-contained. Readers should get a sense of “why” it is an important question to ask and “what” they will learn if they look at it.
    • Deletions: keep in mind that readers only see the simple text. So you may need to add a bit more context into the question to make it apparent to readers “why” this is important.
    • Oversimplifications: explicitly connect to the concept that is being clarified. This can often be done by somehow including the concept in the question.
  • Question scoping: try to phrase the question such that the highlighted piece is the most likely answer, ideally there should be a singular answer.
  • The answer must draw on information from the original.
  • The answer must use plain language that an average adult would understand. You may have to explain difficult terms and concepts.
  • For each QA pair, please check that the answer cannot be found at a later stage in the simplification. The reason is that a simplification may reorder content or drop redundant information to reduce the length. We define information loss globally, so be aware of these cases. Basically, this is a small sanity check that you can do while drafting the QA pair.
  • Avoid: Yes/no questions (rather, reformulate it using a question word: what, how, why, etc.).
  • Avoid: Questions that ask for further simplification or external information (these are not about an information gap).
Figure B.1: Annotation interface for InfoLossQA built with Thresh Heineman et al. (2023).
Appendix C Evaluation Guidelines
We seek to evaluate models that identify information loss caused by simplifying text. These models provide two outputs: (i) a localization of what information was lost, and (ii) a QA pair that elicits the missing information.
C.1 Model Recall of Human-written QA
We aim to determine how many of the human-written QAs are generated by the models. This is a pairwise comparison. You will be given a reference QA and a model QA. Please assign one of the following categories:
  • Fully recalled: the model QA gives the same information as the reference QA
  • Partially recalled: the model QA partially answers the reference QA
  • Not recalled: there is no overlap in the presented content
As our goal is to identify if models identify the same information loss, we should abstract away from surface level dissimilarities and framing of questions. Here are some guidelines to help with this judgment:
  • First, identify the specific unit of information that the reference is asking about by looking both at the reference question, answer and (if needed) the localization. Then check if this information is conveyed by the model answer. The full texts and answer localizations are given for your reference. Note, however, that the model localization may not reflect the generated QA.
  • A useful heuristic is to check how well the reference answer could be replaced by the model-generated answer, even if this may lead to a slightly weird/unintuitive phrasing.
  • Disregard background explanations. When a reference answer explains a technical concept but the model answer does not include this explanation, it can still be fully recalled. We evaluate simplicity separately.
  • Both the reference or the model-generated QA can include more information that cannot be matched to the other QA. For example, the model QA could ask a broader question which entails multiple reference QAs.
  • In examples where both QAs are asking for the same information, but the model answer misinterprets the information, we mark as not recalled.
See the table below for example annotations.
# Reference QA Model QA Rating Comment
1
How were the carbon nanoparticles administered to the participants? This study was a dose-escalation study, meaning that the carbon nanoparticles were administered in increasing doses sequentially to different groups of participants.
What type of study was conducted to test the effects of carbon nanoparticles? An investigator-initiated, randomized controlled, dose-escalation study was conducted.
Full
Disregard background explanation in the original.
2
What condition did the babies in the study have? The babies in the study were diagnosed with moderate-to-severe bronchiolitis.
How was the study conducted? The study was a double-blind, randomized controlled trial on infants (1 to 12 months) who were diagnosed in the emergency department with moderate-to-severe bronchiolitis.
Full
Fully recalled even if phrased differently
3
What statistical tests were used to analyze the results? The results were analyzed using a repeated measures (the procedures were conducted multiple times), mixed design ANOVA test (compares the results from three or more study groups).
What type of statistical analysis was used to analyze the results? A repeated measures, mixed design ANOVA was used to analyze the results.
Full
Disregard background explanation
4
What was done to prevent bias in the trial? The trial was designed as a double-blind trial, meaning neither the patients nor the researchers knew which group of participants had been assigned which combination of sedative drugs.
How was this study conducted? This study was conducted as a clinical trial at a single center. It was double-blinded, meaning that neither the researcher nor the participants knew which treatment each participant received, and it was randomized, meaning the participants were randomly assigned one of the treatments.
Full
Model QA is quite broad but includes the missing bit of information (double-blind).
5
What was the major criterion for patients to participate in the study? The 155 patients who participated in this study had a glaucoma-related diagnosis, meaning they had been diagnosed with an eye condition that causes vision loss.
What kind of health condition did all enrolled patients have? All enrolled participants were diagnosed with a glaucoma-related condition.
Full
Main subject is “major criterion” which the model QA identifies. 155 patients is not the subject of the question, and background explanation “vision loss” can be ignored.
6
How were the carbon nanoparticles administered to the participants? This study was a dose-escalation study, meaning that the carbon nanoparticles were administered in increasing doses sequentially to different groups of participants.
[…]
Partial
The question asks for dose-escalation. The answer shows increasing doses, but it misses calling out “dose escalation.”
7
What specific parts of the digestive system were observed? The study looked at the antrum (part of the upper stomach) and the second duodenal portion (lower part of the stomach).
In this study, where were the muscle contractions assessed? The muscle contractions were assessed in the antrum, which is a part of the stomach.
Partial
8
What other analyses were done of the rsFC strength? Correlation analyses were done to see if the rsFC strength under heroin exposure was connected to the rewarding effect and plasma (in the blood) concentrations of heroin.
Did the researchers also look for a relationship between the strength of the brain connections and the level of morphine in the patients’ blood? Yes, the researchers studied whether there was a correlation between the strength of resting state functional connectivity (rsFC) and the plasma concentration of morphine, a major metabolite of heroin.
Partial
Rewarding effect is missing.
9
What kind of tests were used to assess patients in this study? Participants were assessed with laboratory tests, the United Kingdom screening test […], the Michigan neuropathy screening score […], and the Michigan diabetic neuropathy score […]. In addition, […]
What method was used to assess the level of neuropathy in the patients? The Michigan neuropathy screening score was used to assess the level of neuropathy in all patients.
Partial
One model generates relatively narrow questions which can touch upon a small part of a larger reference question.
10
What kind of factors is the study analyzing to determine how effective the fast track area (FTA) method is? The study is analyzing the impact of a fast tract area (FTA) using both effectiveness measures, such as waiting times and length of stay in the waiting room, and quality measures, such as mortality rates and LWBS rates (patients who leave without being seen by a medical professional).
What is the purpose of the fast track area? The purpose of the fast track area is to reduce waiting times, patient dissatisfaction, and morbidity by providing a separate area for less urgent cases, allowing for faster treatment and better patient flow.
No
Superficially a lot of overlap, but the answer here needs to make clear what is measured not what the fast track area can hypothetically do.
11
What aspects of heart function were increased using the combination of LS and SNP drugs? Left ventricular systolic (the heart’s ability to pump blood) and diastolic (heart relaxation and how long it takes for the ventricles, two chambers of the heart, to fill with blood) functions including LVEF (how much blood is pumped out of the heart each time it contracts (beats)) improved using the combination of LS and SNP drugs.
What was the main conclusion of the study? The study shows that the LS and SNP combination really improved heart function during and after surgery. Using these drugs led to less need for other supporting drugs or equipment. Thus, we suggest using this drug combination before and during surgery.
No
Does not give an answer about the actual heart function.
12
How much did the special questionnaire overestimate calcium intake compared to the 24-hour recall? The questionnaire overestimated the average total calcium intake by 221 mg/d (milligrams per day) compared to the 24-hour recall across racial groups.
What was the variability in daily calcium intake estimated by the 24-hour recall? The variability or standard deviation in daily calcium intake estimated based on the 24-hour dietary recall was 433 milligrams per day.
No
13
How did the other treatments perform in comparison to each other? The results showed that Group 2 that received only the laser treatment showed better results than Group 1 who received only a low-dose of potassium nitrate.
What were the specifics of the treatment groups in the study? The treatment groups in the study were: Group 1 - 5% potassium nitrate, Group 2 - gallium-aluminum-arsenide diode laser with a wavelength of 980 nm, noncontact pulse mode, and power wattage of 0.5 W, Group 3 - combined 5% KNO3 and the diode laser, and Group 4 - placebo (control).
No
Figure C.1: Interface for pairwise evaluation of QA recall. Built with LabelStud.io.
C.2 Quality of Generated QA
Motivation/rationale (all questions).
For each rating, please provide a brief explanation that motivates your choice. For positive ratings, the rationale can be omitted. For negative ratings, explain how the QA pair could be improved to get to a positive rating. This helps us to analyze the answers and to improve the annotation guidelines.
Criterion 1: Question Givenness
A question should be interpretable for a reader. It should only contain concepts (entities, events, or states) that were mentioned in the question context or concepts that are generally known or inferable from mentioned ones. We define question context as follows: For deletions, the context is the entire simple text. For oversimplifications, the context is everything in the simple text up to and including the question localization. For this evaluation, please pretend that you only see the simplified text.
How is the question phrased?
  • $\square$
    Good (reader focused, no new concepts)
  • $\square$
    Bad (e.g., question introduces new concepts, answer leakage, hallucinations)
Criterion 2: Question Localization
When the question seeks more information/clarification about an oversimplified concept, the corresponding span in the simplified text should be highlighted. For “Missing:” highlight the corresponding text in the simple text which discusses the topic in an oversimplified way.
To what extent does the highlight relate to the topic under discussion?
  • $\square$
    Good: the highlight corresponds to the topic that the question discusses
  • $\square$
    Unrelated: the highlighted text does not relate to what the question is asking
  • $\square$
    Missing: there should be a highlight, but there is not (please add highlight…)
  • $\square$
    n/a: the topic under discussion is not part of the simple text (= deletion)
Criterion 3: Answer Simplicity
The answer should be easy to understand. Please focus on the readability and simplicity of the answer. This is different from accuracy which we will evaluate later.
Does the answer contain jargon?
  • $\square$
    The answer is jargon-free
  • $\square$
    The answer contains jargon but it is adequately explained in the answer
  • $\square$
    The answer contains jargon but it is adequately explained in the simplified text
  • $\square$
    The answer contains unexplained jargon
Is the answer standalone?
  • $\square$
    Yes, the answer can be understood without looking at the original
  • $\square$
    No, the answer contains confusing aspects (e.g., unresolved coreferences, abbreviations/acronyms)
Criterion 4: Answerability/Question Relevance
The question should be about an information loss between the original and simplified texts. We evaluate this in two steps: answerability on the original text, and answerability on the simplified text. A question is about an information loss if it is answerable on the original, but unanswerable/only vaguely answerable on the simplified text.
When making your assessment, you may use the answer snippet to help with this evaluation. However, be aware that the snippet may not correctly answer the question. In those cases, disregard the snippet and look for other places which answer the question.
Is the question answerable with the original text?
  • $\square$
    Yes, and there is a single obvious answer
  • $\square$
    Yes, but there could be multiple valid answers
  • $\square$
    No
To what extent is the question answerable with the simplified text (i.e., degree of information loss)? The benchmark for this is the answer on the original text.
  • $\square$
    Fully answerable. Asking it on the simplified text would give the same answer or a closely paraphrased answer as on the original.
  • $\square$
    Partly or vaguely answerable. The simplified text gives some relevant information, but is less specific or exhaustive than the original.
  • $\square$
    Unanswerable.
Criterion 5: Answer Accuracy
The question should be correctly answered. This criterion only applies to questions that can be answered with the original text. Unanswerable questions are out of scope of this investigation. Ideally a question should be specific enough so that there is a singular answer (see Criterion 4: Answerability). If the question is ambiguous/vague (i.e., there are multiple valid answers), its answer has a high chance of being incomplete or only partially answering the question.
Does the provided answer correctly answer the question?
  • $\square$
    Yes
  • $\square$
    Partially, the answer is related but misses information
  • $\square$
    No
Does the provided answer have any hallucinations? Hallucinations are information or claims that cannot be traced back to the original. Disregard general background explanations and elaborations.
  • $\square$
    Good: there are no hallucinations
  • $\square$
    Bad: the answer contains hallucinations
Does the answer snippet correctly answer the question? Imagine it was rephrased into a fluent answer.
  • $\square$
    Yes
  • $\square$
    Partially, the answer snippet is related but misses information
  • $\square$
    No
Figure C.2: Interface for quality assessment of QA (criteria continued in scrollbar). Built with LabelStud.io.